What's the answer for 10 5/8 x 6 2/5
Answer by fion6002(19)

10 5/8 * 6 2/5
= (8*10+5)/8 * (6*5+2)/5 (improper fraction)
= 85/8 * 32/5
= 17/1 * 4/1 (simplified 32 and 8, 85 and 5)
=17*4
=68
or without simplified
= 85/8 * 32/5
= 680 / 40
= 68